OpenStudy (lgbasallote):

The partial pressure of CH4(g) is 0.175 atm and that of O2(g) is 0.250 atm in a mixture of the two gases. a. What is the mole fraction of each gas in the mixture? i know Pt = Pa + Pb and Pa = XaPt do i use those here?

OpenStudy (lgbasallote):

@Kryten do you know this?

OpenStudy (anonymous):

yep you use those equations... so first you sum up the pressures and then divide individual pressures with summed up pressure and you will get percentage of each pressure in mixture

OpenStudy (lgbasallote):

ohhh yeah i see \[X_A = \frac{P_A}{P_T}\]

OpenStudy (lgbasallote):

xa is the mole fraction right?

OpenStudy (anonymous):

yap

OpenStudy (lgbasallote):

lol i was thinking i need \[n_A \; \text{and}\; n_B\] to solve it

OpenStudy (lgbasallote):

what about this.. b. If a mixture occupies a volume of 10.5 L at 65ºC, calculate the number of moles of each gas in the mixture. what do i do then?

OpenStudy (anonymous):

so you go like this: 0,175 + 0,25 = 0,425 and now 0,175 / 0,425 = 0,4117 = 41,17 % of CH4 and 1-0,4117 = 0,5883 = 58,83 % O2

OpenStudy (lgbasallote):

wait...ratio is %?

OpenStudy (lgbasallote):

isnt it like something-over-something?

OpenStudy (anonymous):

no ratio isnt % but that is how you calculate percentage lol

OpenStudy (lgbasallote):

ohhh. so the ratio would just be 0.175/0.425 or 0.4117, etc?

OpenStudy (anonymous):

yep if im not still sleaping that is right :)

OpenStudy (lgbasallote):

so then what about b b. If a mixture occupies a volume of 10.5 L at 65ºC, calculate the number of moles of each gas in the mixture. what do i do there?

OpenStudy (anonymous):

ok so now you need to calculate amount of mixture in moles: pV=nRT--> n= pV/RT = 0,161 mol (watch for units in pV=nRT !!! they must be in K, m3, Pa...) and now you have X= n/n(sum) --> n = X * n(sum) n(CH4)= 0,4117 * 0,161 mol = 0,06628 mol

OpenStudy (lgbasallote):

where did you get 0.161 again? is it given or solved?

OpenStudy (anonymous):

solved from pV=nRT p= 0,425 atm V= 10,5 L T=65 °C

OpenStudy (lgbasallote):

0.425 is the total right?

OpenStudy (anonymous):

yep 0,425 is the total from a. part of the task

OpenStudy (lgbasallote):

why should the units be K, m3 and Pa...shouldnt it be K, L and atm?

OpenStudy (anonymous):

you always need to calculate in standard units! they give you in nonstandard units so you need to recalculate! thats part of the solving! R = 8,314 J mol-1 K-1 so to cross out you need to have temperature in K pressure in Pa and V in m3

OpenStudy (lgbasallote):

wait i thought \[R = 0.0821 \frac{L \cdot atm}{K \cdot mol}\]

OpenStudy (lgbasallote):

those are the standard units i know...

OpenStudy (anonymous):

i dont see what you have written [Math Processing Error][Math Processing Error] is all i see

OpenStudy (lgbasallote):

refresh

OpenStudy (anonymous):

you are from US arent you?! well take a look at this: http://en.wikipedia.org/wiki/International_System_of_Units there you got all standard units but you in US still use nonstandard units like F for temperature etc.

OpenStudy (lgbasallote):

no im not

OpenStudy (anonymous):

ok, calculate how you know just if you get right result... :)

OpenStudy (lgbasallote):

i got 0.1607

OpenStudy (lgbasallote):

so what happened next?

OpenStudy (lgbasallote):

this is the total mole right...so what's the individual mole?

OpenStudy (lgbasallote):

@Kryten ?

OpenStudy (lgbasallote):

never mind i got it